OpenStudy (anonymous):

Which angle corresponds to 1/6 of a circle?

OpenStudy (anonymous):

π4 π6 π3 180°

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

16.6666667% is 1/6 of a circle

OpenStudy (anonymous):

those are pi/4 pi/6 pi/3

OpenStudy (acxbox22):

one full circle is 360 degrees convert into radians and one full circle = 2pi radians

OpenStudy (acxbox22):

divide 2pi by 6 to get 1/6th of a circle in radians

OpenStudy (jdoe0001):

\(\bf \textit{full circle}=2\pi \qquad \cfrac{1}{6}\textit{ of a circle}=\cfrac{1}{6}\cdot 2\pi \to \cfrac{\cancel{2}\pi }{\cancel{6}}\)

OpenStudy (anonymous):

2(3.14)/6?

OpenStudy (jdoe0001):

yes, just simplify it

OpenStudy (anonymous):

you divided 2 and 6

OpenStudy (anonymous):

3pi?

OpenStudy (jdoe0001):

well, yes, common factor, 2 of each

OpenStudy (jdoe0001):

\(\bf \textit{full circle}=2\pi \qquad \cfrac{1}{6}\textit{ of a circle}=\cfrac{1}{6}\cdot 2\pi \to \cfrac{\cancel{2}\pi }{\cancel{6}}\to \cfrac{\pi }{3}\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

gotchA

OpenStudy (jdoe0001):

:)